# GIM: Improved Interpretability for Large Language Models

### v1

@@ 6 Results
Self-repair is prevalent in attention mechanisms: Our analysis reveals that attention self-repair is common across large language models. Figure 2(a) demonstrates this phenomenon by comparing the joint effect of ablating the two largest attention scores simultaneously (y-axis) versus the sum of their individual ablation effects (x-axis). The consistent gap above the diagonal, especially prominent for orange-highlighted instances (top 5% strongest interactions), provides clear empirical evidence of self-repair: individual attention scores show disproportionately small effects when ablated separately. These results validate our theoretical analysis from Section 3.1 in practical model settings. We provide results for other models and datasets in Section 9.9.
TSG attributes more importance to attention scores during attention self-repair: Figure 2(b) compares standard gradient attribution (x-axis) with our Temperature-adjusted Softmax Gradients (y-axis) for the same attention scores. The orange dots represent attention vectors identified as having strong self-repair interactions in the left figure. TSG consistently attributes higher importance to these scores. This shows that TSG successfully addresses the gradient cancellation issue identified in our theoretical analysis by increasing attribution magnitude where standard gradients would otherwise assign low importance due to self-repair. We provide similar results for other models and datasets in Section 9.9.
GIM is more faithful than other gradient-based circuit identification methods: Figure 3 shows comprehensiveness and sufficiency metrics across layers for LLAMA-3.2 1B. GIM significantly outperforms other circuit identification methods in early layers, while Integrated Gradients demonstrates slight advantages in middle and late layers. Similar patterns appear across other models and datasets (Section 9.8).
For all the methods, the scores worsen in the top layers of the LLM. This could either be because they become less faithful or because the upper and lower limits of comprehensiveness and sufficiency change across layers [18]. Since we are interested in the relative differences between the methods, we leave it to future work to understand what causes the drop in the scores in top layers.
GIM is often more faithful than other gradient-based feature attribution methods: Table 1 compares GIM against five other gradient-based attribution methods across multiple datasets and model architectures. GIM drastically outperforms traditional methods like GradientXInput, Integrated Gradients, and DeepLIFT on nearly all evaluations. While AttnLRP [10] outperforms GIM on several dataset-model pairs, GIM achieves the highest comprehensiveness and sufficiency scores for the majority of combinations, establishing it as the most consistently faithful method overall.
Table 1: Comparison of gradient-based feature attribution methods.
@@
All modifications contribute to GIM’s faithfulness improvements: Figure 4 compares how adding different GIM modifications to GradientXInput improves feature attribution faithfulness. Each point represents the average improvement for a dataset compared to the GradientXInput baseline. Blue boxplots show feature attribution performance with LayerNorm freeze + Gradient Normalization, while orange boxplots include TSG as well (full GIM). The results demonstrate that LayerNorm freeze + Gradient Normalization substantially improve both attribution metrics across all models. Adding TSG further enhances attribution quality for most models, with particularly strong improvements for Gemma-2 2B and LLAMA models. However, TSG appears less beneficial for Qwen-2.5 1.5B, where it occasionally reduces attribution faithfulness. Qwen-2.5 1.5B had a hundred-fold larger gradients than the other models, which perhaps is problematic for TSG (see Figure 11). Comprehensive feature attribution results for all modification combinations are available in Table 2.
@@
Figure 4: Relative improvement of adding modifications to GradientXInput. We compare adding layernorm freeze and grad norm with adding also TSG. Each point represents the average statistics of a dataset. The y-axis shows the improvements over GradientXInput.
@@ 7 Discussion
When to amplify signals during self-repair: Our temperature-adjusted softmax gradients (TSG) amplify attribution signals when attention self-repair occurs, but is this amplification justified? We argue that amplification is appropriate for components that were actively contributing to the output before any perturbation, even when self-repair dampens the observed output change. The “firing squad” analogy from Pearl [30] illustrates this well: when two soldiers simultaneously fire at a prisoner (where either shot alone would be fatal), both bear full responsibility despite the redundancy. Even though removing one shooter’s bullet would not change the outcome (a form of self-repair), we should still attribute causal responsibility to both shooters.
Crucially, we should not attribute importance to components that only begin to contribute after perturbation occurs, regardless of whether this redistribution happens immediately (as in attention self-repair) or through backup mechanisms (as in the hydra effect [8]). Consider a historical counterfactual: “What killed Archduke Franz Ferdinand?” If we ablate the assassination in Sarajevo, and in this counterfactual he later dies of cancer, we would not say he died from both being shot and cancer. When a component only contributes after ablation of another component, that compensating component should not receive attribution in the unperturbed state.
@@
Limitations: While our experiments demonstrate that TSG increases both backpropagation signals during self-repair and overall attribution faithfulness, we cannot conclusively establish a causal link between these observations. Temperature adjustment affects backpropagation in multiple ways beyond mitigating self-repair, and these additional effects may contribute to the improved faithfulness we observe.
Furthermore, we did not find an explanation for why TSG improved faithfulness across all models except for Qwen-2.5 1.5B. We observed that its gradient was a hundred-fold larger than the other models’, but could not provide any definite evidence that this caused the decreased faithfulness.
@@ 8 Conclusion
We identified attention self-repair, a phenomenon where softmax normalization redistributes weights during perturbation, causing interpretability methods to underestimate component importance. Our proposed method, Gradient Interaction Modifications (GIM), addresses this through temperature-adjusted softmax gradients (TSG), while also incorporating layer normalization freeze and gradient normalization to handle other issues in transformer interpretability. Our evaluations across multiple models and datasets show that GIM consistently outperforms existing circuit identification and feature attribution methods in faithfulness. By accounting for self-repair during backpropagation, GIM better quantifies the importance of model components, advancing our understanding of neural networks by improving our ability to interpret them.
Acknowledgments and Disclosure of Funding
Thank you, Casper L. Christensen, for the enthusiastic discussions and for proposing and testing ideas for how to improve GIM. Thank you, Jakob Havtorn, Lasse Borgholt, and Søren Hauberg, for discussions and advice.

### v3

@@ 6 Results
Self-repair is prevalent in attention mechanisms: Figure 2(a) compares the effect of ablating the two largest attention scores simultaneously (y-axis) versus the sum of their individual ablation effects (x-axis) for the samples we identified as self-repair. The consistent gap above the diagonal provides clear empirical evidence of self-repair: individual attention scores show disproportionately small effects when ablated separately. These results empirically validate our theoretical analysis from Section 3.1. We provide results for other models and datasets in Section˜A.10. On average, we identified self-repair in the model 65–1200 times per input, depending on the model and dataset.
TSG accurately approximates the joint effect: Figure 2(b) and 2(c) compare standard gradient attribution and TSG with the joint effect. As intended, TSG more accurately approximates the effect of perturbing multiple attention scores simultaneously. We provide similar results for other models and datasets in Section˜A.10.
GIM is more faithful than other gradient-based circuit identification methods: Figure˜3 shows comprehensiveness and sufficiency metrics across layers for LLAMA-3.2 1B. GIM significantly outperforms other circuit identification methods in early layers, while Integrated Gradients demonstrates slight advantages in middle and late layers. Similar patterns appear across other models and datasets (Section˜A.9). Csordás et al. (2025) showed that LLMs primarily use the attention heads in the early layers. Since most of our modifications target the attention heads, this could explain the pattern. Across all methods, the scores worsen in the top layers of the LLM. This could either be because they become less faithful or because the upper and lower limits of comprehensiveness and sufficiency change across layers Edin et al. (2025). In Section˜A.1, we show that GIM outperforms the other circuit discovery methods on the Mechanistic Interpretability Benchmark.
GIM is often more faithful than other gradient-based feature attribution methods: Table˜1 compares GIM against five other gradient-based attribution methods across multiple datasets and model architectures. GIM significantly outperforms traditional methods, such as GradientXInput, Integrated Gradients, and DeepLIFT, in nearly all cases. While AttnLRP outperforms GIM on several dataset-model pairs, GIM achieves the highest comprehensiveness and sufficiency scores for the majority of combinations, establishing it as the most consistently faithful method overall.
Table 1: Comparison of the faithfulness of gradient-based feature attribution methods. The best scores for each model-dataset pair are shown in bold.
@@
All modifications contribute to GIM’s faithfulness improvements: Figure˜4 compares how adding different GIM modifications to GradientXInput improves feature attribution faithfulness. Each point represents the average improvement for a dataset compared to the GradientXInput baseline. Blue boxplots show feature attribution performance with LayerNorm freeze + Gradient Normalization, while orange boxplots include TSG as well (full GIM). The results demonstrate that LayerNorm freeze + Gradient Normalization substantially improves both attribution metrics across all models. Adding TSG further enhances attribution quality for most models, with particularly strong improvements for Gemma-2 2B and LLAMA models. However, TSG appears less beneficial for Qwen-2.5 1.5B, where it occasionally reduces attribution faithfulness. Qwen-2.5 1.5B had a hundred-fold larger gradients than the other models, which perhaps is problematic for TSG (see Figure˜11). While this inter-model variance is undesirable, it is even larger for established methods such as IG and GxI. Comprehensive feature attribution results for all modification combinations are available in Table˜4.
@@
Figure 4: Relative improvement of adding modifications to GradientXInput. We compare adding layernorm freeze and grad norm with also adding TSG. Each point represents the average statistics of a dataset. The y-axis shows the improvements over GradientXInput.
@@ 7 Discussion
Should gradients estimate the joint effect? We designed TSG to approximate the effect of perturbing multiple attention scores simultaneously to address the self-repair problem. However, is this always desirable? Whether to approximate the effect of perturbing one or multiple attention scores depends on the causal question being asked. If a perturbation would cause changes in multiple interacting attention scores, then we should estimate the joint effect; otherwise, we should estimate the individual effect. However, when perturbing early-layer variables, determining which attention scores in deep layers are affected becomes difficult because the number of potential causal paths grows exponentially with depth. We naively assumed that perturbations would always impact all interacting attention scores. This assumption may fail for certain inputs, models, and types of perturbations, which could explain why GIM underperforms other methods for Qwen 1.5B in Table˜1 and in the late layers in Figure˜3. TSG is a naive solution to the self-repair problem; future work should investigate methods to estimate whether a perturbation will cause changes in multiple interacting attention scores, enabling a more principled selection between joint and individual effect estimation.
@@
Limitations: While our experiments demonstrate that TSG improves the faithfulness of the attributions and accurately approximates the effect of joint ablation during self-repair, we cannot conclusively establish a causal link between these observations. Temperature adjustment affects backpropagation in multiple ways beyond mitigating self-repair, and these additional effects may contribute to the improved faithfulness we observe. Furthermore, we did not find an explanation for why TSG improved faithfulness across all models except for Qwen-2.5 1.5B. We observed that its gradient was a hundred-fold larger than the other models’, but could not provide any definite evidence that this caused the decreased faithfulness.
@@ 8 Conclusion
We identified attention self-repair, a phenomenon where softmax normalization redistributes weights during perturbation, causing interpretability methods to underestimate component importance. Our proposed method, GIM, addresses this through TSG, while also incorporating layer normalization freeze and gradient normalization to handle other issues in transformer interpretability. Our evaluations across multiple models and datasets show that GIM consistently outperforms existing circuit identification and feature attribution methods in faithfulness. By accounting for self-repair during backpropagation, GIM more accurately quantifies the importance of model components, thereby advancing our understanding of neural networks and enhancing our ability to interpret them.
9 Reproducibility statement
Our code is available at https://github.com/JoakimEdin/gim. In this code repository, we provide commands for running every experiment. In addition to the code, we provide experimental details in Section˜5.2. We also provide information about the required licenses for using the datasets in Section˜A.4, and the computational resources we used for the experiments in Section˜A.4.
Acknowledgement